OpenStudy (anonymous):

F(x) = 2x^(3) + 3x^(2) + 7x + 4, a = 4. Find (f^(-1))'(a).

OpenStudy (anonymous):

this is just like what we worked in prev question

OpenStudy (anonymous):

I don't know if I did this correctly, I took the easy route of doing it. I graphed the actual equation, and found where '4' is in the Y axis table, and then took the X point and considered it the answer.

OpenStudy (anonymous):

I tried finding the inverse of it, but it was too complicated because I ended up getting stuck with many terms with Y's and it was hard to simplify

OpenStudy (anonymous):

no need to find inverse do u know the formula for deriative of inverse function?

OpenStudy (anonymous):

Hmm, no. I registered for the class late and missed the first day.

OpenStudy (anonymous):

http://oregonstate.edu/instruct/mth251/cq/Stage6/Lesson/inverseDeriv.html \[\frac{d}{dx} f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

suppose f^(-1)(4)=b then\[\frac{d}{dx} f^{-1}(4)=\frac{1}{6b^2+6b+7}\]

OpenStudy (anonymous):

so u just need to evaluate b

OpenStudy (anonymous):

So what you are saying is that I just need to put the equation in the denominator and plug in the a value?

OpenStudy (anonymous):

f'(x)=6x^2+6x+7 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

If I plug in 4 into the equation you posted, I would get 1/127

OpenStudy (anonymous):

b is not 4 ........... b is f^(-1)(4)